### v2

# ZClassifier: Temperature Tuning and Manifold Approximation via KL Divergence on Logit Space

Abstract
We present ZClassifier, a classifier model that replaces deterministic logits with Gaussian-distributed latent variables, trained via KL-regularized variational sampling. By explicitly regularizing the latent logits toward a standard Gaussian distribution, ZClassifier improves both classification calibration and out-of-distribution (OOD) detection.
We evaluate the model on CIFAR-10 as the in-distribution dataset, and SVHN, Gaussian noise, and Uniform noise as OOD inputs. ResNet-based ZClassifier achieves near-perfect OOD detection (AUROC = 0.9994, AUPR = 0.9994, FPR@95 = 0.0000) while maintaining 99% OOD classification accuracy. VGG-based ZClassifier, in contrast, shows weaker latent separation (AUROC = 0.8333, FPR@95 = 0.4114). When KL regularization is removed, the model fails to detect OOD entirely (AUROC = 0.0122), indicating latent space collapse.
These findings support two hypotheses: (1) Gaussian latent modeling of logits introduces a useful inductive bias for robust uncertainty estimation Lee et al. (2018); Venkataramanan et al. (2023); Kim and Pavlovic (2014), and (2) skip connections in ResNet architectures improve the expressiveness and regularization of latent variables under variational objectives Dieng et al. (2019); Nagayasu and Watanabe (2023).
ZClassifier thus offers a practical and theoretically grounded solution for uncertainty-aware image classification with strong OOD robustness.
1 Introduction
Deep neural networks have achieved strong performance in supervised classification across visual domains. Yet, their reliance on deterministic logit vectors passed through a softmax activation often results in overconfident predictions, particularly on inputs far from the training distribution Guo et al. (2017). Such overconfidence hinders the reliable deployment of neural classifiers in risk-sensitive settings.
Moreover, the standard softmax classifier encodes no explicit inductive bias about class-wise uncertainty or latent structure. While recent efforts in contrastive and probabilistic representation learning have emphasized the importance of structured latent geometry Khosla et al. (2020); Oh et al. (2020), conventional classifiers lack mechanisms to reflect class-conditional manifolds or variance.
To address these shortcomings, we introduce ZClassifier, a classification framework in which each class logit is modeled as a Gaussian latent variable. Given an input, the model produces a distribution over logits—parameterized by class-specific means and variances—and aggregates them by sampling and averaging. A KL divergence term regularizes the predicted logit distribution toward a standard normal prior, thereby promoting well-behaved latent geometry and uncertainty calibration.
This design provides two key advantages:
  1. 1.
    Uncertainty-aware prediction: Gaussian latent modeling enables structured confidence that separates in-distribution (InD) and out-of-distribution (OOD) inputs without requiring post-hoc calibration.
  2. 2.
    Latent expressiveness via skip connections: When instantiated with ResNet backbones, the model benefits from skip connections that prevent latent collapse and encourage diverse class representations even under KL regularization.
We empirically validate ZClassifier on CIFAR-10 as the in-distribution dataset, and evaluate OOD detection using SVHN, Gaussian noise, and Uniform noise. Our results show that the ResNet-based ZClassifier achieves near-perfect separation of OOD samples (AUROC = 0.9994, FPR@95 = 0.0000), while the VGG-based variant exhibits significantly degraded performance. An ablation without KL regularization fails entirely at OOD detection, highlighting the importance of latent distribution control.
ZClassifier offers a simple yet effective probabilistic extension of standard discriminative classifiers. It builds upon insights from variational inference Kingma and Welling (2014) and recent advances in Gaussian-based uncertainty modeling Lee et al. (2018); Venkataramanan et al. (2023), while exploiting the structural advantages of residual architectures Dieng et al. (2019); Nagayasu and Watanabe (2023).
2 Related Work
Calibration and Softmax Limitations. The softmax classifier with deterministic logits is known to produce poorly calibrated confidence estimates, especially on out-of-distribution (OOD) inputs Guo et al. (2017). Temperature scaling addresses this by rescaling logits:
$\hat{p}(y\mid x)=\mathrm{softmax}\left(\frac{f_{\theta}(x)}{T}\right),$
where $T>0$ is optimized post-hoc on a validation set. Although this improves calibration, it does not introduce any uncertainty modeling or distributional structure into the logits themselves.
More principled approaches include Monte Carlo Dropout Gal and Ghahramani (2016) and Deep Ensembles Lakshminarayanan et al. (2017), which estimate uncertainty via multiple forward passes. However, they do not modify the logit layer itself, and their computational cost scales linearly with ensemble size or dropout samples.
Latent Structure and Regularization. Several methods introduce structure into intermediate embeddings $h(x)$ via contrastive or metric learning. Supervised contrastive learning Khosla et al. (2020) encourages class-wise compactness in $h$, but the classifier head $g_{\phi}$ remains linear and deterministic. Hedged Instance Embedding Oh et al. (2020) augments embeddings with Gaussian noise $h\sim\mathcal{N}(\mu,\Sigma)$ to model aleatoric uncertainty, leading to KL-regularized metric objectives.
ZClassifier departs from these approaches by applying stochastic modeling directly at the logit level. This imposes structure on the class-wise decision surface rather than only on the feature representation.
Gaussian Logit Modeling and Variational Inference. Our method is most closely related to probabilistic classifiers that model logits as random variables. The Mahalanobis classifier of Lee et al. (2018) assumes that class-conditional features follow Gaussian distributions and scores test inputs by their Mahalanobis distance to class means. MAPLE Venkataramanan et al. (2023) extends this by learning class-specific latent Gaussians and calibrating uncertainty through distance-based OOD scoring.
ZClassifier adopts a similar Gaussian assumption, but imposes it per-logit rather than per-feature. Each class logit is modeled as $z_{k}\sim\mathcal{N}(\mu_{k}(x),\sigma_{k}^{2}(x))$, and the mean over samples is used for prediction. A KL regularization term enforces alignment with $\mathcal{N}(0,1)$ priors across logits, inspired by the variational inference framework Kingma and Welling (2014); Dhuliawala et al. (2024). This helps prevent overconfident predictions and collapses in the latent structure.
Latent Collapse and the Role of Skip Connections. Recent works have examined how skip connections affect the informativeness of latent variables. Dieng et al. (2019) propose Skip-VAE, showing that feeding latent variables into intermediate generative layers prevents posterior collapse and improves latent usage. Similarly, Nagayasu and Watanabe (2023) analyze residual networks from a Bayesian perspective and show that skip connections preserve generalization by stabilizing the free energy bound under depth growth.
We observe similar benefits in ZClassifier: ResNet-based models, equipped with skip connections, consistently learn better-separated and more expressive latent distributions under KL regularization compared to VGG-based counterparts. This provides architectural evidence for the interaction between network topology and latent uncertainty modeling.
3 Methodology
3.1 Gaussian Logit Modeling
Probabilistic Logit Distributions. Instead of producing a fixed logit vector, ZClassifier models each class logit as a Gaussian-distributed latent variable. For input $x^{(i)}$, the network outputs class-wise mean $\mu^{(i)}\in\mathbb{R}^{K}$ and variance $\sigma^{(i)2}\in\mathbb{R}^{K}$. This yields a factorized distribution over logits:
$q_{i}(z)=\prod_{c=1}^{K}\mathcal{N}(z_{c}\mid\mu^{(i)}_{c},\sigma^{(i)2}_{c}).$
During inference, prediction is made via $\arg\max_{c}\mu^{(i)}_{c}$, but training uses full distributions for uncertainty-aware loss.
Class-Conditional Prototypes. We define the target distribution for label $y_{i}=c$ as:
$p_{y_{i}}(z)=\mathcal{N}(\mu^{*}_{c},I),$
where $\mu^{*}_{c}$ is a one-hot vector: 1 at index $c$, 0 elsewhere. This anchors each class in logit space such that the true class is centered at 1 and all others at 0.
KL Regularization. The KL divergence between predicted logits and the class prototype is:
$\mathcal{L}_{\mathrm{KL}}^{(i)}=\frac{1}{2}\sum_{c=1}^{K}\left[(\mu_{c}^{(i)}-\mu^{*}_{y_{i},c})^{2}+\sigma_{c}^{(i)2}-1-\log\sigma_{c}^{(i)2}\right].$
This regularizes the predicted distribution toward the structured Gaussian prior.
Total Loss. The model is trained using a weighted sum of cross-entropy and KL loss:
$\mathcal{L}^{(i)}=-\log\frac{\exp(\mu^{(i)}_{y_{i}})}{\sum_{c}\exp(\mu^{(i)}_{c})}+\lambda\cdot\mathcal{L}_{\mathrm{KL}}^{(i)}.$
We set $\lambda=10$ by empirical tunings.
Reparameterization and Latent Dimensionality. To support stochastic training, we sample $z_{c}=\mu_{c}+\sigma_{c}\cdot\epsilon$ using $\epsilon\sim\mathcal{N}(0,1)$. Each class may have latent dimension $d>1$, in which case we average across dimensions:
$\bar{z}_{c}^{(i)}=\frac{1}{d}\sum_{j=1}^{d}z_{c,j}^{(i)}\sim\mathcal{N}\left(\mu_{c}^{(i)},\frac{\sigma_{c}^{(i)2}}{d}\right).$
This reduces the variance of the final prediction, ensuring stability even when predicted variances are large.
3.2 Model Variants and Baselines
ZClassifier (ours). Our primary model applies the above Gaussian logit formulation using a feature extractor backbone (ResNet18 or VGG11), followed by a shared fully connected head to produce $\mu$ and $\log\sigma^{2}$. Latent sampling and KL divergence are applied per instance, and the model is trained end-to-end.
No-KL Ablation. To test the role of KL regularization, we train the same architecture but set $\lambda=0$, removing the prior matching constraint. This allows the model to optimize purely for classification, potentially leading to degenerate latent distributions and failure to detect OOD samples.
Softmax Classifier. As a baseline, we implement a standard discriminative model using the same feature extractor and a linear head to output logits directly. This model is trained with cross-entropy only and does not model uncertainty or apply latent regularization.
3.3 Benefits of Gaussian Logit Modeling
Adaptive Calibration. Each predicted variance $\sigma_{c}^{2}$ acts as an instance-specific temperature controller, enabling the model to learn when to be confident or uncertain. Unlike static temperature scaling Guo et al. (2017), this is learned per class and per sample.
Uncertainty-Aware OOD Detection. High KL divergence or inflated predicted variances often indicate mismatch from the prototype, suggesting OOD input. We exploit this for effective threshold-based detection.
Geometric Regularization. By constraining logits toward $\mathcal{N}(1,1)$ for true classes and $\mathcal{N}(0,1)$ for others, the model avoids uncontrolled logit scaling and learns shape-sensitive, scale-invariant representations.
Efficiency. Unlike ensemble or MC-based methods, ZClassifier supports single forward-pass uncertainty estimation and performs robustly without additional inference overhead.
4 Experiments
This section presents an empirical evaluation of ZClassifier under classification, uncertainty estimation, and distributional robustness settings. We compare four model variants: KL-regularized ZClassifier and standard softmax classifiers, each instantiated with ResNet-18 (Model A) and VGG-11 (Model B) backbones.
Our evaluation covers: (1) classification performance across classes via precision-recall analysis, (2) robustness to Gaussian input noise, (3) out-of-distribution (OOD) detection using KL divergence scoring under SVHN, Gaussian, and Uniform shifts, and (4) latent space structure via covariance matrices and PCA.
These experiments reveal the impact of architectural choice and KL regularization on predictive reliability, uncertainty calibration, and latent geometry.
4.1 Classification Performance Comparison
Table 1: Test Classification Report for ZClassifier and Softmax/No-KL Baselines. Best values in each column are bolded.
Class ZClassifier Softmax / No-KL
Prec-A Prec-B Rec-A Rec-B Prec-A Prec-B Rec-A Rec-B
0 0.73 0.84 0.94 0.83 0.85 0.89 0.83 0.85
1 0.93 0.89 0.93 0.93 0.73 0.84 0.98 0.95
2 0.90 0.76 0.68 0.71 0.79 0.80 0.79 0.75
3 0.78 0.63 0.61 0.69 0.70 0.65 0.68 0.76
4 0.79 0.81 0.87 0.74 0.76 0.74 0.86 0.86
5 0.71 0.74 0.84 0.71 0.69 0.88 0.83 0.64
6 0.82 0.85 0.94 0.84 0.91 0.82 0.82 0.92
7 0.96 0.81 0.81 0.87 0.95 0.95 0.78 0.81
8 0.92 0.90 0.90 0.90 0.95 0.94 0.86 0.91
9 0.95 0.91 0.87 0.90 0.96 0.91 0.72 0.88
Accuracy 0.84 0.81 0.81 0.83
Analysis.
ResNet-based ZClassifier achieves the highest classification accuracy (84%) with strong precision-recall balance, particularly in ambiguous or underrepresented classes such as 0, 3, and 5. It benefits from structured latent representations induced by KL-regularized Gaussian logits.
VGG-based ZClassifier shows weaker recall in several classes (e.g., 3, 5) despite reasonable precision, indicating less expressive latent geometry. This is consistent with prior findings that deeper but skip-less networks may struggle with feature diversity.
Softmax classifiers perform comparably in top-1 accuracy, but exhibit high recall and low precision in some classes (e.g., class 1), suggesting overconfident predictions. The No-KL variant achieves competitive accuracy (83%) but suffers from unstable variance and inconsistent calibration—highlighting the importance of KL regularization in latent structure enforcement.
In summary, ZClassifier with ResNet backbone best balances accuracy, class sensitivity, and stability—showcasing the synergy between latent Gaussian modeling and architectural expressiveness.
4.2 Latent Structure Analysis
To understand how architectural choices and logit modeling influence latent structure, we analyze the class-wise covariance, correlation, and geometric alignment of predicted logits. We compare KL-regularized ZClassifier models to softmax-based baselines across both ResNet and VGG architectures.
Figure 1: Latent structure visualizations for ZClassifier. Top row: class-wise covariance and correlation matrices for ResNet (left) and VGG (right). Middle: Frobenius norm difference of covariance (left) and correlation (right). Bottom-left: PCA projection of logits (ResNet: blue, VGG: orange). Bottom-right: Bartlett’s $p$-values per dimension ($\alpha=0.05$).
Covariance and Correlation Matrices.
Figure 1 (top row) presents covariance and correlation matrices of class logits from ResNet (Model A) and VGG (Model B) within the ZClassifier framework. ResNet reveals diverse inter-class dependencies and structured off-diagonal patterns. In contrast, VGG exhibits a flatter, more isotropic covariance profile, indicating collapsed or entangled class representations.
The Frobenius norm of the covariance difference is 55.63, suggesting substantial architectural influence. The correlation matrices differ with a norm of 1.33—less than for raw covariance but still notable. This supports the claim that skip connections in ResNet help preserve structured, class-dependent uncertainty even under KL regularization.
PCA Visualization.
The PCA projection (Figure 1, bottom-left) shows that ResNet logits (blue) cluster by class more tightly and distinctly than VGG logits (orange), which are more diffuse and linearly distributed. This reflects a more disentangled latent structure in ResNet-based ZClassifier.
Bartlett’s Test for Variance Homogeneity.
We compute Bartlett’s test per logit dimension (Figure 1, bottom-right) to assess the homogeneity of logit variance across classes. Multiple dimensions fall below the $\alpha=0.05$ threshold, especially in ResNet, confirming that ZClassifier learns class-specific distributions with distinct variances—consistent with our Gaussian modeling objective.
Figure 2: Latent structure visualizations for softmax-based classifiers. Top: class-wise covariance and correlation matrices. Middle: Frobenius norm differences (Cov and Corr). Bottom-left: PCA projection. Bottom-right: Bartlett’s $p$-values. Note the degenerate geometry compared to Figure 1.
Softmax Comparison.
Figure 2 shows the same metrics for softmax classifiers. Both ResNet and VGG variants exhibit minimal covariance diversity. The Frobenius norm difference between their covariance matrices is only 0.17, indicating near-identical structure. PCA projections form narrow manifolds with no visible class clustering, and Bartlett’s $p$-values are close to zero for all dimensions, suggesting collapsed and homoscedastic logits.
Summary.
Compared to softmax models, KL-regularized ZClassifier—with ResNet backbone—yields a more informative and diverse latent space. This validates our architectural choice and objective formulation: KL divergence regularization, combined with skip-connected feature extractors, enhances class separability and geometric richness in the logit space.
4.3 Calibration Robustness under Logit Perturbation
To evaluate the stability and calibration robustness of each model, we inject Gaussian noise into the predicted logits at inference time and measure classification accuracy. This simulates logit-level uncertainty and tests how well each model maintains discriminative power when confidence is perturbed.
Figure 3: Accuracy under additive Gaussian logit noise. ResNet ZClassifier remains stable up to STD $\sim$1.0, while Softmax and NoKL variants degrade rapidly.
Table 2: Calibration Accuracy by Logit Noise STD
Noise STD ResNet ZC VGG ZC Softmax NoKL ZC
0.00 0.8408 0.8122 0.8148 0.8326
0.22 0.8391 0.8071 0.7975 0.8157
0.44 0.8331 0.7932 0.7459 0.7563
0.67 0.8259 0.7647 0.6685 0.6917
0.89 0.7988 0.7180 0.5973 0.6066
1.11 0.7513 0.6631 0.5324 0.5398
1.33 0.6808 0.6114 0.4648 0.4725
1.56 0.6126 0.5544 0.4046 0.4276
1.78 0.5467 0.5002 0.3751 0.3902
2.00 0.4970 0.4530 0.3379 0.3549
Observations.
Figure 3 and Table 2 reveal stark differences in robustness:
  • ResNet ZClassifier maintains high accuracy up to noise STD 1.0, degrading gradually. Its performance remains above 50% even at STD 2.0, demonstrating strong resilience to uncertainty.
  • VGG ZClassifier performs reasonably up to STD 0.9, but falls off more steeply than ResNet. This reflects its more fragile latent geometry, as seen in previous sections.
  • Softmax Classifier shows rapid decay after STD 0.5. This is expected, as deterministic logits have no internal regularization or uncertainty modeling to compensate for perturbation.
  • NoKL ZClassifier shows similar degradation to Softmax, confirming that without KL regularization, latent variance is unstructured and does not provide robustness.
Interpretation.
These results suggest that KL-regularized Gaussian modeling of logits plays a critical role in calibration robustness. The structured latent distributions learned by ZClassifier act as an inductive prior that resists small perturbations, unlike softmax logits, which are scale-sensitive and unbounded. The advantage is magnified when combined with skip-connected architectures like ResNet.
4.4 Latent Logit Geometry on Validation Set
To further understand the internal geometry learned by each model, we visualize the latent logit vectors of validation samples from CIFAR-10 using Principal Component Analysis (PCA). We project the high-dimensional class logits into 2D space and color each point by ground-truth class.
(a) ResNet ZClassifier
(b) VGG ZClassifier
(c) Softmax Classifier
(d) No-KL ZClassifier
Figure 4: PCA projection of latent logit vectors for each model on CIFAR-10 validation set. Each point represents a sample and is colored by its class label.
ResNet ZClassifier.
Figure 4(a) shows tight, radially organized clusters aligned with class identity. Classes are well-separated in logit space and exhibit angular structure, suggesting directional encoding consistent with the KL-regularized Gaussian latent objective.
VGG ZClassifier.
Figure 4(b) demonstrates a similar geometric “fan-shaped” structure, but with more class overlap. Despite the Gaussian loss, VGG fails to fully disentangle logit space. This aligns with the hypothesis that skip connections aid in preserving class-conditional directions.
Softmax Classifier.
In Figure 4(c), logits form a dense, entangled cloud with partial class clusters. Without any latent regularization, softmax logits collapse toward magnitude-based separation. There is no explicit variance modeling, so the geometry lacks class-wise anisotropy.
No-KL ZClassifier.
Figure 4(d) reveals scattered and overlapping clusters with poor separation. Despite the stochastic output, the lack of KL alignment results in unstructured logit space. Classes bleed into one another, highlighting the role of KL loss in shaping semantic directions.
Summary.
These visualizations confirm that: (1) ZClassifier learns more structured latent spaces than softmax-based baselines, (2) KL regularization is essential for geometric disentanglement, and (3) residual architectures like ResNet amplify class separation under Gaussian latent assumptions.
4.5 Linear Discriminability via Fisher LDA
To complement PCA-based geometric analysis, we apply Fisher’s Linear Discriminant Analysis (LDA) to evaluate class-wise linear separability in the latent logit space. LDA projects class logits onto a lower-dimensional space that maximizes inter-class variance while minimizing intra-class spread. This is a useful proxy for testing how well the model organizes its representation under supervised guidance.
(a) ResNet ZClassifier
(b) VGG ZClassifier
(c) Softmax Classifier
(d) No-KL ZClassifier
Figure 5: Fisher LDA projections of latent logit vectors for four model variants. Each dot corresponds to a validation sample, colored by true class. Clear separation indicates strong class encoding.
ResNet ZClassifier.
In Figure 5(a), class clusters are distinctly separated and linearly aligned. Each class forms a dense, low-variance region, with inter-class gaps reflecting high Fisher scores. This indicates that KL-regularized latent logits align not only with class means but also with class directions in a discriminative space.
VGG ZClassifier.
In contrast, Figure 5(b) shows partial class overlap and fuzzier boundaries. While the general alignment is preserved, classes are less compact, particularly for class 3 and 5. This aligns with previous PCA findings that VGG’s latent structure is flatter and less anisotropic.
Softmax Classifier.
Figure 5(c) shows tightly packed class clusters with significant entanglement. Without stochasticity or regularization, the logits collapse into a small region, reducing linear separability. Class 7, 8, and 9 show substantial overlap, indicating degraded discriminability under linear projections.
No-KL ZClassifier.
Figure 5(d) reveals high overlap and diffuse boundaries. Although stochastic, the lack of KL regularization fails to shape the latent space. Class directions are unstable, and class variances are unregulated, confirming the structural collapse observed in other analyses.
Summary.
Fisher LDA projections confirm that ZClassifier—especially with ResNet backbone—produces logit spaces that are both semantically structured and linearly discriminable. KL divergence not only regularizes variance but also implicitly enforces class-specific directions, facilitating downstream calibration and OOD detection tasks.
4.6 Nonlinear Structure via t-SNE Projection
To visualize global and local semantic structure of the latent logit space, we apply t-distributed Stochastic Neighbor Embedding (t-SNE) to the predicted logits of each model. While PCA and LDA reveal linear separability, t-SNE captures nonlinear manifolds and local neighborhood fidelity across classes.
(a) ResNet ZClassifier
(b) VGG ZClassifier
(c) Softmax Classifier
(d) No-KL ZClassifier
Figure 6: t-SNE embeddings of latent logit vectors on CIFAR-10 validation set. Colors indicate ground-truth class. Separation, density, and compactness of clusters reflect learned latent geometry.
ResNet ZClassifier.
In Figure 6(a), the logit embeddings form highly structured, radially symmetric clusters with large inter-cluster margins. Each class is clearly defined with tight local grouping, indicating strong preservation of class manifolds in high dimensions. This reflects the effect of KL-regularized Gaussian sampling combined with skip-connected feature extractors.
VGG ZClassifier.
Figure 6(b) shows more fragmented clusters. While most classes remain separated, there are multiple split clusters per class and greater local noise. This suggests that VGG lacks sufficient expressiveness to shape logits as compact distributions in t-SNE space, despite having the same Gaussian loss.
Softmax Classifier.
Figure 6(c) reveals spherical but entangled class regions. Although global structure is evident, local boundaries are less distinct. This is expected as softmax logits are not regularized and only reflect hard decision surfaces rather than probabilistic embeddings.
No-KL ZClassifier.
In Figure 6(d), we observe loosely defined clusters with significant overlap. Interestingly, the stochasticity from Gaussian sampling alone produces some variance, but in the absence of KL regularization, class directions and cluster centers are unaligned—leading to unstable latent encoding.
Summary.
t-SNE results corroborate previous findings: ZClassifier with ResNet backbone induces both global separability and local cohesion in latent logit space. In contrast, removing KL or using less expressive architectures leads to more entangled and diffuse embeddings, even under identical task supervision.
4.7 Gaussianity of Logits via GMM Ellipse Analysis
To evaluate whether the predicted class logits align with class-conditional Gaussian structures, we fit a Gaussian Mixture Model (GMM) to the PCA-projected logit space and visualize per-class covariance ellipses. This provides a geometric test of whether the latent structure exhibits class-specific normality, as expected under our KL-regularized objective.
(a) ResNet ZClassifier
(b) VGG ZClassifier
(c) Softmax Classifier
(d) No-KL ZClassifier
Figure 7: 2D GMM ellipses over PCA-projected logit space. Ellipses represent class-wise Gaussian components (means and covariances). Top and right margins show marginal KDEs along each principal axis.
ResNet ZClassifier.
As shown in Figure 7(a), GMM ellipses for each class are compact, distinct, and oriented in semantically aligned directions. This indicates that the model learns near-isotropic Gaussians with low overlap, consistent with its KL-regularized latent objective. Marginal KDEs show unimodal, well-separated peaks along both principal axes.
VGG ZClassifier.
Figure 7(b) shows elliptical structures with higher anisotropy and partial overlap. Some ellipses are skewed or elongated, suggesting instability in class-wise variance modeling. While still Gaussian-like, the lack of skip connections results in less regularized latent clusters.
Softmax Classifier.
Figure 7(c) reveals spherical but overlapping ellipses. Despite forming separable clusters in PCA, the underlying distributions lack variance regularization, leading to uncontrolled scaling and mixture ambiguity. Marginal distributions are broad and multi-modal, confirming the absence of latent distributional control.
No-KL ZClassifier.
Figure 7(d) illustrates the least structured geometry. Ellipses are large, overlapping, and misaligned, with high inter-class confusion. Although the model samples from Gaussians, the absence of KL divergence means these samples do not converge to meaningful class prototypes. KDEs reflect blurred, entangled distributions.
Summary.
GMM ellipse analysis confirms that KL-regularized ZClassifier (especially with ResNet) best satisfies the latent Gaussian prior assumption. This leads to more coherent uncertainty modeling and downstream improvements in OOD detection and calibration stability.
4.8 Out-of-Distribution Detection with KL Divergence
To assess the robustness of ZClassifier to out-of-distribution (OOD) inputs, we measure the KL divergence between each predicted Gaussian logit distribution and a uniform target distribution. A high KL divergence indicates strong deviation from the learned in-distribution patterns. Figure 8 shows the KL score histograms for in-distribution CIFAR-10 and three types of OOD inputs: SVHN, Gaussian noise, and Uniform noise.
(a) ResNet-18 on SVHN
(b) ResNet-18 on Gaussian
(c) ResNet-18 on Uniform
(d) VGG-11 on SVHN
(e) VGG-11 on Gaussian
(f) VGG-11 on Uniform
(g) NoKL Variant on SVHN
(h) NoKL Variant on Gaussian
(i) NoKL Variant on Uniform
Figure 8: Histograms of KL divergence scores between predicted logits and a uniform distribution. Each plot compares in-distribution CIFAR-10 samples and OOD samples under various ZClassifier variants. Clear separation is achieved with KL-regularized Gaussian modeling.
Quantitative Evaluation.
We report standard OOD detection metrics—AUROC, AUPR, and FPR@95—in Table 3. These metrics evaluate the model’s ability to distinguish between in-distribution (CIFAR-10) and OOD samples (SVHN, Gaussian noise, Uniform noise) based on KL divergence scores computed from the predicted Gaussian logits.
The ResNet-based ZClassifier demonstrates exceptional discriminative capability across all three OOD types. It achieves an AUROC of at least 0.999, an AUPR above 0.9994, and a false positive rate at 95% recall (FPR@95) effectively zero. This indicates that the KL divergence values for in- and out-distribution samples are almost completely separable. We attribute this performance to two factors: (1) the expressiveness of the ResNet backbone, which captures a structured logit geometry; and (2) the KL regularization term, which explicitly enforces parametric uncertainty and separation across class-conditional distributions.
In contrast, the NoKL variant of the ZClassifier collapses in all OOD scenarios. Despite having the same backbone and Gaussian logit structure, the absence of the KL term leads to degenerate behavior. The AUROC falls below 0.03, and the FPR@95 reaches 1.0000—a complete failure in distinguishing OOD inputs. This suggests that the model no longer encodes meaningful class uncertainty in its latent space. Instead, logits become overconfident and collapse toward uninformative directions, resulting in overlapping distributions for in- and out-of-distribution data.
The VGG-based ZClassifier exhibits intermediate performance. On Gaussian and Uniform noise, the model attains AUROC scores of 0.95 and 0.93, respectively, which are competitive. However, the model degrades significantly on the SVHN dataset (AUROC: 0.83, AUPR: 0.56, FPR@95: 0.41). This result may reflect two aspects: the VGG backbone’s weaker inductive bias compared to ResNet and a latent space geometry that is less robust to natural image shifts. The SVHN dataset, although semantically different from CIFAR-10, contains well-structured visual content that is more difficult to reject via density-only cues, unlike pure noise.
In summary, the results highlight that (1) KL regularization is essential for OOD separability in Gaussian logit models, (2) ResNet-18 backbones induce more reliable uncertainty-aware decision boundaries, and (3) model expressiveness and regularization jointly determine the fidelity of probabilistic latent structure necessary for robust detection.
Table 3: OOD Detection Performance via KL Divergence (Higher AUROC/AUPR, Lower FPR@95 are better).
Model OOD Type AUROC AUPR FPR@95
ResNet SVHN 0.9994 0.9994 0.0000
ResNet Gaussian Noise 0.9997 0.9998 0.0000
ResNet Uniform Noise 0.9992 0.9996 0.0000
VGG SVHN 0.8333 0.5573 0.4114
VGG Gaussian Noise 0.9528 0.9511 0.2152
VGG Uniform Noise 0.9298 0.9455 0.4424
NoKL SVHN 0.0122 0.1540 1.0000
NoKL Gaussian Noise 0.0208 0.3076 1.0000
NoKL Uniform Noise 0.0184 0.3093 1.0000
Classification Stability.
We also evaluate OOD detection from a binary classification perspective: given a test sample, the task is to determine whether it originates from the in-distribution (CIFAR-10) or from one of the OOD sources (SVHN, Gaussian noise, Uniform noise), using KL divergence as the decision score. We apply a fixed threshold—optimized on a held-out subset—for each model.
The ResNet-based ZClassifier consistently achieves high binary detection accuracy across all OOD types:
  • SVHN: 98.6% total accuracy (InD recall = 0.95, OOD recall = 1.00)
  • Gaussian noise: 97.0% total accuracy
  • Uniform noise: 97.0% total accuracy
These results indicate that the model maintains stable decision boundaries in the presence of both natural and synthetic domain shifts. Notably, the predicted KL divergence values for OOD samples are sharply concentrated in high-score regions, facilitating confident rejection without degrading InD classification.
By contrast, the NoKL variant of the ZClassifier collapses under the same evaluation. Regardless of the OOD type, the model fails to distinguish samples, resulting in binary classification accuracy near or below 50%—essentially chance level. In particular:
  • All FPR@95 metrics for NoKL are 1.0000, showing no ability to reject OOD
  • Precision and recall for OOD class are near zero in all reports
This reinforces the idea that the KL divergence term is not merely auxiliary—it is structurally essential for forming separable and meaningful latent distributions under uncertainty.
Implications.
The strong separation of KL score distributions between InD and OOD samples validates our core hypothesis: KL divergence from class prototypes acts as a statistically grounded, discriminative signal for distributional shift.
By modeling each class logit as a Gaussian and enforcing alignment with isotropic class prototypes via KL regularization, the ZClassifier creates a probabilistic latent space in which:
  1. 1.
    In-distribution samples exhibit low KL values due to close alignment with learned class priors.
  2. 2.
    OOD samples yield high KL scores due to misaligned means and inflated variances, even when feature-level similarity is nontrivial (e.g., SVHN digits).
This behavior emerges directly from the model’s forward pass—no auxiliary classifier, energy function, or external uncertainty estimator is required. The KL divergence is a by-product of the model’s generative assumption and can be computed efficiently at inference.
Ultimately, ZClassifier serves as both a classifier and an uncertainty estimator in a unified framework, offering principled, calibrated decision-making under distributional shift.
5 Conclusion
We proposed ZClassifier, a probabilistic classifier that replaces deterministic logits with diagonal Gaussian latent variables, trained via KL-regularized variational sampling. This model unifies uncertainty calibration, logit normalization, and latent structure modeling under a single objective. Unlike softmax-based discriminative models, ZClassifier directly parameterizes per-class distributions, enabling robust detection of out-of-distribution (OOD) samples via KL divergence.
Empirical results demonstrate that our approach achieves:
  • Near-perfect OOD detection across natural (SVHN) and synthetic (Gaussian, Uniform noise) domains with AUROC $\geq$ 0.999 using only KL scores,
  • Improved calibration robustness under logit-level perturbations without requiring post-hoc temperature scaling,
  • Superior latent geometric structure, as confirmed via PCA, LDA, t-SNE, and GMM ellipses—especially when paired with expressive backbones like ResNet.
Ablation studies further highlight that KL regularization is essential to prevent latent collapse, and that architectural choices (e.g., skip connections in ResNet) interact meaningfully with the Gaussian logit assumption to produce linearly and nonlinearly separable latent spaces.
In conclusion, ZClassifier offers a simple yet powerful alternative to ensemble-based uncertainty estimation or generative modeling. It achieves calibrated, uncertainty-aware predictions and reliable OOD detection in a single forward pass, using only a lightweight modification to standard classifier architectures. We believe this framework serves as a general-purpose template for probabilistic classification in real-world, shift-prone settings.
6 Discussion
6.1 Limitations
Our evaluation of ZClassifier has been limited to CIFAR-10, a small-scale dataset with only 10 flat classes. This leaves open questions regarding the model’s scalability to more complex domains. For instance, in tasks with a large number of classes (e.g., CIFAR-100 or ImageNet), the diagonal Gaussian assumption for each class may become insufficient to preserve separation. Likewise, the model has not yet been tested in structured prediction tasks such as semantic segmentation, where each pixel is associated with spatially correlated class predictions. In these cases, the independence assumption between class logits may no longer hold.
Another limitation is ZClassifier’s untested performance in class-imbalanced or long-tailed distributions. While Gaussian modeling may intuitively capture uncertainty for low-frequency classes, there is no guarantee that KL divergence will regularize minority logits effectively without explicit balancing. Finally, the assumption that each class logit distribution is isotropic Gaussian—centered at either $\mathcal{N}(1,1)$ (true class) or $\mathcal{N}(0,1)$ (false class)—may break down in fine-grained or hierarchically structured settings, where feature overlap and semantic correlation between classes are significant Ridnik et al. (2023); Hsu et al. (2019).
6.2 Future Work
Hierarchical Gaussian Logit Modeling.
Future work could adopt tree-structured or hyperbolic latent priors Nickel and Kiela (2017); Liu and Nickel (2020), enabling classes to share covariance structure based on semantic hierarchy. For example, a parent-child Gaussian prior could allow fine-grained subclasses to inherit from superclass prototypes, improving expressiveness and calibration in taxonomic datasets.
Imbalanced and Few-Shot Learning.
ZClassifier could be extended to handle long-tailed distributions via KL loss reweighting Cao et al. (2019) or variance-scaling for minority classes Ren et al. (2020). In few-shot regimes, recent work has shown that probabilistic embeddings improve generalization Snell et al. (2017); Zhang et al. (2020), suggesting that ZClassifier may offer strong sample efficiency through its uncertainty-aware latent space.
Integration with Foundation Models.
Replacing the softmax head in ViT Dosovitskiy et al. (2021) or CLIP Radford et al. (2021) with a Gaussian logit head could yield more calibrated large-scale classifiers. Techniques like zero-shot classifier tuning or vision-language pretraining could be combined with our logit modeling to enhance robustness and interpretability Liu et al. (2023); Zhang et al. (2023).
Diffusion Guidance.
Classifier-guided diffusion sampling Dhariwal and Nichol (2021); Ho and Salimans (2022) is a promising application. ZClassifier can provide both mean-based guidance and variance-aware modulation for more controllable generation. This fits well into emerging frameworks that mix probabilistic and generative signals for improved fidelity and safety.
In sum, future research will aim to generalize ZClassifier’s probabilistic structure to hierarchical, low-data, or open-world scenarios while preserving the interpretability and sampling efficiency enabled by its Gaussian foundation.